40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1669     Entry time: Thu Jun 11 22:14:10 2009
Author: Alberto 
Type: Update 
Category: LSC 
Subject: DD Handoff for the Short DOFs completed 

This afternoon I tuned the handoff script for the SRC, after that Rob eralier during the day had already adjusted that for PRC. To do that, I followed the procedure in the Wiki.

  • I measured the OL transfer function of the single demod path and of the double demod path and tuned thier gains so that they matched
  • I tuned the double demod pahses of PD_7 and PD_8 in order to reduce the offset in the PD_x_I signals

After that the SRC could get locked with the double demod signals. the open loop transfer function emasurement on the PRC loop showed that it was nearly unstable. Rob reduced a little its gain to improve the stability.

The DD handoff is now working and we can get back to locking the interferometer.

ELOG V3.1.3-